### v2

# Anyon braiding and telegraph noise in a graphene interferometer

Abstract
The search for anyons, quasiparticles with fractional charge and exotic exchange statistics, has inspired decades of condensed matter research. Quantum Hall interferometers enable direct observation of the anyon braiding phase via discrete interference phase jumps when the number of encircled localized quasiparticles changes. Here, we observe this braiding phase in both the $\nu=1/3$ and $4/3$ fractional quantum Hall states by probing three-state random telegraph noise (RTN) in real-time. We find that the observed RTN stems from anyon quasiparticle number $n$ fluctuations and reconstruct three Aharonov-Bohm oscillation signals phase shifted by $32\pi/$ 3, corresponding to the three possible interference branches from braiding around $n$ (mod 3) anyons. Hence, we reveal the anyon braiding phase using new methods that can readily extend to non-abelian states.
Introduction
The fractional quantum Hall (FQH) effects, where electrons are confined to two spatial dimensions and exposed to large magnetic fields, have long been predicted to host emergent fractionally charged excitations that obey neither fermionic nor bosonic exchange statistics [1, 2, 3, 4, 5]. These quasiparticles fall into two classes, abelian and non-abelian anyons, associated with FQH states at different filling factor $\nu$. For abelian anyons, exchanging two quasiparticles (i.e. braiding) results in the many-body wavefunction acquiring a complex phase, while in the non-abelian case, the wavefunction evolves by a unitary transformation, enabling fault-tolerant topological quantum computation that has yet to be experimentally realized [6, 7, 8, 9, 10].
FQH Fabry-Pérot (FP) interferometers enable direct measurements of the anyon braiding phase. [11, 12]. By partially backscattering current at two quantum point contact (QPC) constrictions, the conductance through the FP cavity modulates with the phase accumulated by edge-traveling quantum Hall (QH) quasiparticles [5, 13]. When the number of localized cavity quasiparticles encircled by the interfering edge changes, the resulting phase shift in the interference signal is equivalent to twice the fundamental exchange phase (modulo $2\pi$) [14, 15]. However, since the observed discrete phase jumps represent charging events in the cavity, resultant Coulomb effects on the interference path can pose complications in probing the braiding effect. In particular, the Coulomb coupling between the bulk and edge must be accounted for when extracting the anyon braiding phase from a small number of phase jump events [16]. Consequently, mitigating Coulomb coupling while maintaining electrostatic tunability has been a major challenge in semiconductor-based interferometers [17, 15].
Figure 1: Graphene FQH interferometer telegraph noise. (A) False-color scanning electron microscope image of a representative device depicting an interference measurement with a single edge channel. Scale bar: $1\mu$m. We measure the conductance $G=I_{d}/(V_{D}^{+}-V_{D}^{-})$ where $I_{d}$ is the measured drain current and $V_{D}^{\pm}$ are the measured voltages at the indicated locations. The top graphite gate is etched into 8 separately tunable regions controlled by voltages $V_{LG}$, $V_{LSG}$, $V_{PG}$, $V_{MG}$, $V_{RSG}$, and $V_{RG}$. Additionally, voltages $V_{LBG/RBG}$ applied to suspended bridges above the two QPCs independently set their transmissions. See Fig. S1 for more details. (B) Hall conductivity $\sigma_{xy}$ and longitudinal conductivity $\sigma_{xx}$ measured with contacts in the left reservoir of the device highlighting the robust plateaus at $\nu=0$, $1/3$, $1$, and $4/3$, which are the quantum Hall states we use in this work. (C) Derivative of $\sigma_{xy}$ showing the expected magnetic field dependence of the well-developed quantum Hall states. (D) Conductance $G$ across the left/right QPC tuned by $V_{LBG/RBG}$ with the right/left QPC deactivated and the bulk in $\nu=4/3$. See Fig. S2 for more details on QPC tunings. Inset: schematics of the edge channel configuration and tunneling at a QPC. As $V_{LBG/RBG}$ increases, the QPC opens more and transmits the outer integer edge (pink) first and eventually transmits the inner fractional edge (blue) with weak backscattering. (E) Three-state RTN in the conductance $G$ when the fractional edge of $\nu=4/3$ is weakly backscattering (corresponding to the tunings marked by the black/purple dots for the left/right QPC in D). Top inset: cartoons of the three possible interference branches from braiding around $n$ (mod 3) anyons, corresponding to the three conductance levels. Right inset: histogram of the conductance over time with 1000 bins. (F) Extracted sinusoidal oscillations from histograms over approximately 8 minutes for each $V_{PG}$ voltage. Each data point shows the central conductance value of a gaussian fit to a histogram peak, shaded by the total counts summed under the gaussian. See Fig. S3 for similar data in $\nu=1/3$. All data is at $T=20$ mK and $B=12$ T unless otherwise noted.
Graphene-based interferometers offer several advantages for realizing anyon braiding, both for the abelian and non-abelian cases. First, both QPCs [18, 19] and interferometers [20, 21, 22, 23, 24] exhibit robust electrostatic tunability for integer QH states. Second, graphite gates reduce Coulomb coupling to reveal Aharonov-Bohm (AB) oscillations [21], which previously required complex quantum well structures in GaAs [17, 15]. Simultaneously, these atomically-flat gates enhance FQH states by screening out charge disorder [25, 26] while allowing precise control of QPC transmission [21, 27, 28] and FP cavity filling factor [29]. Encouragingly, bilayer graphene has displayed a multitude of even-denominator FQH states which may host non-abelian anyons with relatively large energy gaps [30, 31, 32, 33, 34], as required for topologically-protected qubits. Less explored even-denominator states have also been observed in monolayer graphene [35, 36].
In this work, we report direct observation of abelian anyon braiding using high-visibility Aharonov-Bohm interference in single atomic layer graphene. Owing to the comprehensive electrostatic control of the device, we are able to tune to both $\nu=1/3$ and $\nu=4/3$ at a fixed magnetic field. Strikingly, we observe three-level random telegraph noise (RTN) in each of these FQH states, which was predicted [37, 38, 39] but never observed in previous experiments [15, 16, 17, 40, 41, 42]. Counterintuitively, we show how fluctuations in anyon number within the cavity enable us to directly observe the anyonic braiding phase with minimal complications due to Coulomb effects.
Constructing a graphene fractional quantum Hall interferometer
We construct our device by performing a series of previously described nanolithography steps [21, 29] on a monolayer graphene-based van der Waals heterostructure encapsulated with insulating hexagonal boron nitride [43] and conducting graphite gates. By gating with separated top graphite regions, we define the interferometer cavity and construct two QPCs that introduce backscattering between the opposite chirality QH edge channels, as shown in Fig. 1A. The quality of the graphene channel is maintained in part due to the encapsulating graphite gates, as evident by well-developed FQH states (Fig. 1B,C) at $T=20$ mK. Unlike previous FQH interferometers in GaAs, we can tune the electron density in-situ in graphene to access many FQH states while holding the magnetic field fixed. In this work, we operate at $B=12$ T (unless noted otherwise) and change density to tune between $\nu=4/3$, $\nu=1$, and $\nu=1/3$, enabling a direct comparison of interference between each state.
We begin by tuning the device such that the interferometer cavity and adjoining contact regions are set to $\nu=4/3$ by applying appropriate voltages $V_{MG}$, $V_{LG}$, and $V_{RG}$, while we define the cavity boundary by setting the plunger gate ($V_{PG}$) and QPC split-gate regions ($V_{LSG/RSG}$) to lie within $\nu=0$. We tune two additional metallic bridge gates over the QPCs (applying $V_{LBG/RBG}$) to independently set the QPC transmissions while keeping all other gates fixed. The conductance across each QPC transitions from near 0 to a plateau at $e^{2}/h$, and finally to near $(4/3)e^{2}/h$ with increasing $V_{LBG/RBG}$ as shown in Fig. 1D. The resonances in the conductance in these scans indicate a relatively soft confining potential at this QPC tuning [27, 44]. Other tunings at $\nu=1/3$ and $\nu=1$ have fewer resonances, though we find that the general behavior of the interference signal is unaffected by these transmission resonances. It is then straightforward to tune each QPC to weakly backscatter the fractional edge channel of $\nu=4/3$ by setting $V_{LBG/RBG}$ as indicated in Fig. 1D.
Three-state anyon telegraph noise
By measuring the conductance through the interferometer with the fractional edge of $\nu=4/3$ weakly backscattering at both QPCs, we discover sporadic switches between three discrete levels as a function of time i.e. three-state RTN (Fig. 1E). From a histogram showing how these levels are weighted over a 50 minute scan, we extract the average conductance and the total weight in time associated with each level. Then, we step the plunger gate $V_{PG}$ to modify the area [21, 17, 45, 46], or equivalently the total charge [29, 47] contained in the FP cavity, and repeat the process. Plotting the resulting average conductance values shaded by their weight in time as a function of $V_{PG}$ reveals three interwoven sinusoidal oscillations separated in phase by $32\pi/$ 3 (Fig. 1F).
Each one of these sinusoids, separated in phase by $32\pi/$ 3, represent one of the three $n$ (mod 3) possible phase “branches” associated with braiding the interfering edge quasiparticles around a system composed of $n$ $e/3$ abelian anyons [37, 38, 39].
Demonstrating exchange statistics with RTN is a departure from the typical method used in previous interferometer studies, in which discrete phase jumps in the continuously modulating AB phase $\theta_{AB}(B,V_{PG})$ were argued to represent the addition or subtraction of a quasiparticle in the interferometer bulk [15, 16, 40]. These studies rely on modulating the AB phase using magnetic field $B$ and plunger gate $V_{PG}$ to probe irregularly spaced phase jumps due to anyon localization. In order to extract the anyon braiding phase, however, the experimentally observed phase jumps need to be analyzed with consideration given to the electrostatic coupling between the QH edge and localized bulk states. These analyses often become complicated as the bulk-edge capacitive coupling, the position of the chemical potential relative to the bulk Landau levels, and interferometer area $A$ can drastically impact interference behavior, such as pivoting the interference signal into the well-studied “Coulomb-dominated” regime [12, 16, 45, 48, 49]. It has also been shown that integer quantum Hall fillings can produce similar fractional phase jumps due to the Coulomb coupling effect described above [24, 29, 50, 51].
The measurement scheme presented in our work, i.e. observing a time-dependent interference phase $\phi$ due to the RTN, offers a key benefit in this regard. Here the parameters $(B,V_{PG})$ used to modulate the AB phase
$\theta_{AB}(B,V_{PG})=2\pi\frac{e^{*}}{e}\frac{A(V_{PG})B}{\phi_{0}}$
can be held fixed while the quasiparticle number $n(t)$ fluctuates in time. Then, the expression for the total interference phase may be simplified as
$\phi=\theta_{AB}(B,V_{PG})+n(B,V_{PG},t)\ \theta_{a}\rightarrow\theta_{AB}+n(t)\ \theta_{a}$ (1)
where $\theta_{a}=2\pi/3$ is the braiding phase for an $e^{*}=e/3$ anyon around a localized counterpart (assuming that Coulomb interactions are well screened) and $\phi_{0}=h/e$. Considering $2\pi$ periodicity of $\phi$, the resulting braiding phase can only be one of three values varying by $32\pi/$ 3, reflecting the total number of localized anyons $n(t)$ (mod 3) at a given time. The device accordingly remains static in configurable parameter space $(B,V_{PG})$ while observing conductance switching from real-time fluctuations $n(t)$. Therefore, as this method demonstrates the existence of all three branches associated with the exchange statistics of charge $e/3$ anyons, we construct a complete representation of the state’s braiding outcomes near a fixed device configuration set by $B$, $V_{PG}$, and density (tuned via $V_{MG}$). Additionally, by sampling a large number of phase jumps while clearly resolving three branches at each $V_{PG}$ tuning, we demonstrate that Coulomb interactions are indeed negligible in our devices, allowing observation of $\theta_{AB}$ from the evolution of the individual branches shown in Fig. 1F at fixed $B$.
Aharonov-Bohm magnetic field dependence
Figure 2: Aharonov-Bohm magnetic field trend of $\bm{\nu=1/3}$ branches. (A) 100 repeated $V_{PG}$ sweeps in $\nu=1$, where each sweep takes 35 s, demonstrating the temporal stability of oscillations in the integer state. (B) Magnetic field dependence of the oscillations in $\nu=1$ demonstrating Aharonov-Bohm periodicity since the magnetic field period yields $\phi_{0}/\Delta B_{1}=0.80\ \mu\text{m}^{2}$, consistent with lateral depletion tending to shrink the area enclosed by the edge state bound by a lithographic design of $1.1\ \mu\text{m}^{2}$. (C) 100 repeated $V_{PG}$ sweeps in $\nu=1/3$, the same sweeping parameters as in (A), demonstrating the stochastic fluctuations induced by the RTN that could easily be mistaken for structureless noise, especially if averaged over by a slow sweep. (D) 6 of the $V_{PG}$ sweeps plotted to demonstrate switching between the three sinusoidal branches within a single scan. (E) 2D histograms showing the probability of measuring a given conductance for each $V_{PG}$ value and its trend with the magnetic field. All three branches (partially highlighted) are made evident by this plot, and each trends identically with a flux super-period in the magnetic field corresponding to $e/3$ anyons.
Figure 3: Telegraph noise dependence on filling in $\bm{\nu=1/3}$. (A) Conductance versus time taken at different values for $V_{MG}$, indicated by color (traces are arbitrarily vertically shifted by ordering of $V_{MG}$). The 300 seconds shown are from longer 50 minute datasets. (B) Hall data taken from the left half of the device with density tuned via $V_{LG}$. The color shade bar indicates the associated range shown for $V_{MG}$ in (C-E) below. Color scale on $R_{xx}$ spans from 20 mK (blue) to 600 mK (red). (C) Average branch switching rate along the plateau. Error bars indicate uncertainty associated with the switching identification algorithm and background noise (see SM section 2). (D) Standard deviation of conductance computed for $G(t)$ at fixed $V_{MG}$. (E) The spectral densities of conductance plotted on a log-log scale (traces are arbitrarily vertically shifted by ordering of $V_{MG}$). Dotted lines are fit to the power law dependence $f^{\alpha}$ in the low frequency regime. Inset cartoons indicate the anyonic quasiparticle density of states versus energy, with the vertical dashed lines indicating the Fermi level for a given filling. (F) The exponent $\alpha$ extracted from the fits of the power spectrum within the frequency range associated with the RTN.
Figure 4: Temperature dependent telegraph noise and visibility in $\bm{\nu=1/3}$. (A) Conductance versus time taken at different temperatures $T$ indicated by color (traces are arbitrarily vertically shifted by ordering of $T$). The 600 seconds shown are subsets of longer 50 minute datasets used in subsequent analysis. (B) Oscillation visibility extracted from $V_{PG}$ sweep measurements (21 repeated scans averaged per point). The exponential decay fit to the oscillation visibility results in a characteristic temperature $T_{0}=110$ mK. (C) Low-temperature saturation and high-temperature activation of the average switching rate $\tau_{s}^{-1}$ observed for two fillings: at the center and at the edge of the $\nu=1/3$ plateau. Shown fit lines follow the empirical formula of Eqn. 2.
Measuring how the magnetic field $B$ changes the total interference phase is crucial in determining if the interferometer falls into the AB or Coulomb-dominated regime [12, 5]. We start for comparison with the $\nu=1$ integer QH state. By repeatedly sweeping the plunger gate, we observe the expected single-sinusoidal oscillations devoid of phase jumps (Fig. 2A). As we change $B$, we observe magnetic-field dependent AB interference (Fig. 2B), where the magnetic field period $\Delta B_{1}$ yields an area $\phi_{0}/\Delta B_{1}=0.80\ \mu\text{m}^{2}$, as expected for interference of electrons in $\nu=1$. A similar measurement in $\nu=1/3$ with varying $B$ and $V_{PG}$ produces an irregular 2D map (Fig. 2C) since repeated plunger gate sweeps are not consistent due to the frequent RTN fluctuations. However, we observe that each plunger gate scan contains stochastic switching between 3 sinusoidal branches, again shifted by $32\pi/$ 3 (Fig. 2D). By superimposing 100 of these sweeps and plotting the histogram of total recorded conductance values over $V_{PG}$, we are able to sample enough instances of the AB oscillation signal of each branch to recreate the three interwoven sinusoids seen in the previous measurement technique discussed in Fig. 1. Fig. 2E exhibits the three interwoven branches at different magnetic fields. As the magnetic field decreases (from the top to bottom panels), we observe that each branch continuously drifts in the increasing $V_{PG}$ direction, creating a negative slope of constant phase (dashed line) analogous to the integer case of Fig. 2B. For $\nu=1/3$, the corresponding flux super-periodicity $\Delta B_{1/3}$ (the magnetic field period for a single branch to return to its initial $V_{PG}$ coordinate) yields $3\phi_{0}/\Delta B_{1/3}=0.83\ \mu\text{m}^{2}$. This value is what one expects for interfering $e/3$ quasiparticles such that $\phi_{0}\rightarrow h/(e/3)=3\phi_{0}$. We similarly observe AB interference and a flux super-periodicity on the fractional edge of $\nu=4/3$ (Fig. S4), while the integer edge shows typical electron AB interference with no observable RTN.
The fact that all tested configurations show AB interference (i.e. a negative slope of constant phase with respect to $V_{PG}$ and $B$) instead of Coulomb-dominated effects indicates that our graphene interferometer is sufficiently screened by the graphite gates to suppress long-range Coulomb interaction, as required for observing braiding signatures [12, 15, 16]. Furthermore, the observed $3\phi_{0}$ magnetic-flux super-period in both $\nu=1/3$ and $\nu=4/3$ agrees with the $3\phi_{0}$ super-period observed near the center of the $\nu=1/3$ plateau in GaAs-based FP interferometers [15, 16], suggesting that the quasiparticle gap is large enough to prevent continuous addition of quasiparticles [52].
Telegraph noise across the quantum Hall plateau
We gain insight into the nature of the quasiparticle fluctuations by tracking how the RTN evolves with density (filling) across the FQH plateau. Fig. 3A exhibits subsets of 50 minute measurements of the conductance, showing the $V_{MG}$ dependent RTN across the $\nu=1/3$ plateau. The trace colors indicate the select values of $V_{MG}$ that correspond to the subrange of the $\nu=1/3$ QH plateau as shown in Fig. 3B. Of immediate note are clear changes in both the switching rate and amplitude as $V_{MG}$ steps across the plateau. The average switching rate from the complete 50 minute traces, $\tau_{s}^{-1}$, shown in Fig. 3C, exhibits steep rises near the edges of the plateau and a relatively constant switching rate in between. Fig. 3D shows the standard deviation, $\sigma_{G}$, of the diagonal conductance $G(t)$, providing a direct measure of switching amplitude.
The striking behavior of the anyon branch switching rate and amplitude along the plateau reveals that the quasiparticle fluctuations and coherence are closely correlated with the compressibility of the FQH state. A natural explanation would be to account for the increase in switching rate with an increase in the number of localized states, and thus anyon trapping sites, near the edges of the FQH plateau. However, a simple increase in the individual switching rates of a constant number of traps across the plateau cannot be ruled out with just this observation alone.
The noise spectral density of the diagonal conductance, $S_{G}$, can provide additional information supporting the hypothesis of an increasing trap number near the edge of the plateau. Fig. 3E shows $S_{G}$ as a function of frequency $f$, obtained from the Fourier transform of $G(t)$ at various $V_{MG}$ values across the plateau region. We find that $S_{G}(f)$ follows a power law scaling behavior ${\sim}f^{\alpha}$ in the low frequency regime corresponding to the RTN signal. Fig. 3F exhibits the scaling exponent $\alpha$ obtained from the line fits shown in Fig. 3E. In the middle of the plateau, the exponent $\alpha$ remains close to -2, which is expected from RTN with a single switching timescale [53]. However, near the edges $\alpha$ sharply increases and deviates away from this stable value. It is known that when many instances of RTN with different switching timescales are convolved together, $S_{G}$ is expected to approach $1/f$, or $\alpha=-1$ [54]. Accordingly, the change of $\alpha$ near the edges of the plateau suggests that there are more unique switching timescales at the edge of the plateau than in the middle. This observation, together with the switching rate behavior discussed above, supports that the increasing branch switching incidents across the FQH plateau are indeed associated with increasing localized anyon states in the bulk.
Temperature scaling
We also study how the RTN changes as a function of temperature $T$. Fig. 4A shows $G(t)$ near the middle of the $\nu=1/3$ plateau at $T$ ranging from 20 mK to 220 mK. As temperature increases, the switching amplitude decreases corresponding to a decrease in AB oscillation visibility. Fig. 4B shows separately measured AB oscillation visibility as a function of $T$ using plunger gate sweep measurements (similar to Fig. 2C-E). The visibility decreases following an exponential decay as $T$ increases, with a characteristic energy scale $T_{0}=110$ mK similar to that observed in $\nu=1/3$ in GaAs [15].
In addition to the decrease in visibility, the RTN also shows an increase in switching rate as temperature increases. Fig. 4C shows temperature dependent $\tau_{s}^{-1}$ for two representative data sets, one at the center (red) and the other at the edge (yellow) of the plateau. While switching rate indeed increases quickly at the high temperature limit, there is a dramatic flat saturation of $\tau_{s}^{-1}$ for both data sets at low temperatures, below about $100$ mK and $50$ mK for the plateau center and edge, respectively. We find that an empirical formula which combines a term motivated by an Efros-Shlovskii (ES) form of variable-range hopping (VRH) [55, 56] with an additional temperature-independent transition rate, $\tau_{s,0}^{-1}$,
$\tau_{s}^{-1}=\tau_{s,0}^{-1}+\tau_{s,1}^{-1}(T)\ e^{-(T_{ES}/T)^{1/2}}$ (2)
can provide a reasonable fit to the data as shown by the dashed lines in Fig. 4C. Here $\tau_{s,1}^{-1}(T)\sim 1/T$ and $T_{ES}$ is the characteristic ES VRH temperature scale. While VRH has been used in the past to model longitudinal conduction due to localized states in the quantum Hall effect [56, 57, 58, 59], at the present moment the physical mechanism behind the temperature scaling of $\tau_{s}^{-1}$ here remains nebulous. We comment that there is a wealth of additional information present in the RTN signal, including how the probabilistic weightings of the phase branches (see the right inset of Fig. 1E) evolve with $T$ and $V_{MG}$, that could be analyzed in future studies to inform a microscopic model of the trapping sites present in FQH states.
While we were able to provide evidence that the anyon fluctuations are closely associated with the number of anyon localized states, and that they have consistent temperature scaling properties, it remains unclear whether they arise from device-specific considerations or rather from the fundamental nature of FQH states in monolayer graphene. Based on the uniformity of the AB-modulating sinusoids in Figs. 1F and 2E, we postulate that the anyons are neither hopping from nor getting trapped within regions close to the QPCs, as even small changes in environmental charge would drastically change their tunings and the resultant visibility of the interference signal. Additionally, although fluctuations have not been reported in the same manner in GaAs heterostructure-based fractional interferometry experiments [40, 16, 15] and more recently in bilayer graphene [60], we note that the presence of slower or faster switching timescales could make initial observations and interpretation of anyon fluctuations challenging with standard DC measurement techniques.
Outlook
In this work, we have shown that anyonic quasiparticle fluctuations and the resulting RTN in a FP interferometer can be used as a powerful tool to exhibit the complete set of braiding phases and exchange statistics present in abelian states. These techniques pave the way for future interferometer measurements of non-abelian anyon braiding in even-denominator FQH states, where the small energy gap may enhance the RTN and limit the range of AB phase modulation. Several theoretical studies have already discussed which signatures would be expected in such a case where RTN is present for non-abelian states [37, 38, 39]. Moreover, our ability to observe the slow equilibration of anyons enables future experiments to probe nontrivial anyon dynamics and to dynamically control anyon number in similar devices. This is particularly relevant to the ultimate goal of constructing a topological qubit with non-abelian anyons, which requires an understanding of how to selectively control and modulate the quasiparticle number [61, 62, 63].
Note added: During the preparation of this manuscript, we became aware of a concurrent work using a similar graphene device [64].
Acknowledgments
We thank Abhishek Banerjee, Yuval Ronen, and Steven H. Simon for helpful discussions. The major part of the experiment was supported by DOE (DE-SC0012260). J.R.E. acknowledges support from ARO MURI (N00014-21-1-2537) for sample preparation, measurement, characterization, and analysis. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 20H00354 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. M.E.W. and A.Y. acknowledge support from Quantum Science Center (QSC), a National Quantum Information Science Research Center of the U.S. Department of Energy (DOE). B.I.H. acknowledges support from NSF grant DMR-1231319. Nanofabrication was performed at the Center for Nanoscale Systems at Harvard, supported in part by an NSF NNIN award ECS-00335765.
Author contributions: T.W., J.R.E., and P.K. conceived of and designed the experiment. T.W. and D.H.N. created the van der Waals heterostructure. T.W. performed the nanofabrication. T.W. and J.R.E. performed the measurements. M.E.W. and A.Y. provided the measurement cryostat and collaborated on discussions and analysis. K.W. and T.T. provided the hexagonal boron nitride crystals. T.W., J.R.E., and P.K. analyzed the data and wrote the manuscript with contributions from B.I.H.